### v2

# Continuous symmetry entails the Jordan algebra structure of quantum theory

Abstract
Symmetry postulates play a crucial role in various approaches to reconstruct quantum theory from a few basic principles. Discrete and continuous symmetries are under consideration. The continuous case better matches the physical needs for mathematical models of dynamical processes and is studied here. Applying the representation theory of the orthomodular lattices, we show that the continuous symmetry, together with three further requirements, entails that the underlying mathematical structure of a finite-dimensional generalized probabilistic theory becomes a simple Euclidean Jordan algebra. The further requirements are: spectrality, a strong state space and a condition called gbit property.
Keywords: foundations and reconstructions of quantum theory; quantum information; Euclidean Jordan algebras; quantum logics
1 Introduction
The generalized probabilistic theories (GPTs [3, 5, 6, 17]) are commonly used as generic models to gain a better understanding of the probabilistic and information theoretic foundations of quantum physics and quantum computing or to reconstruct quantum theory from a few basic principles. In the various reconstructions [17], symmetry postulates play a crucial role. In many cases [3, 15, 18], the symmetries under consideration are discrete. Continuous symmetries would better match the physical needs for mathematical models of dynamical processes and are considered in other approaches [11].
So do we here and study the impact of the continuous symmetry on the underlying mathematical structure of the finite-dimensional GPTs. We show that this becomes a simple (irreducible) Euclidean Jordan algebra, when the continuous symmetry, spectrality, a strong state space and one further condition are given. Our proof is based on the representation theory of the orthomodular lattices [14, 22]. This requires the covering property, which we shall derive from the more information theoretical gbit-property - our last condition.
Spectral order unit spaces become our generic mathematical structure and are outlined in section 2. First findings are presented in section 3. The gbit-property is introduced and studied in section 4. Continuous symmetry and the main result are finally elaborated in section 5.
2 The generic framework
In quantum theory, self-adjoint linear operators represent the observables. The usual operator product does not remain self-adjoint, only the Jordan product does, but the physical meaning of the Jordan product of non-commuting observables is not clear. Therefore the question arises what would become an appropriate mathematical structure for the system of observables, when we do not insist on the existence of a product from the beginning.
[…]
$\left\|a\right\|=inf\left\{s\in\mathbb{R}|-s\mathbb{I}\leq a\leq s\mathbb{I}\right\}$
for $a\in A$.
In quantum theory, the projection operators $p^{*}=p=p^{2}$ form the so-called quantum logic [13, 14, 22, 25, 26] and play a special role in the spectral resolutions of the observables. The spectral resolution is an important feature of the mathematical formalism of quantum theory and is needed to define the potential numerical measurement outcomes of an observable (its spectrum) and the probability distributions over them. Without the product we cannot define the projections via $p^{2}=p$. However, there is another characterization of the projections in the self-adjoint part of a von Neumann algebra $M$; the projections are the extreme points of the convex set $\left\{a\in M|0\leq a\leq\mathbb{I}\right\}$ [2]. This characterization remains available in any order unit space $A$ and we consider
$L_{A}:=ext\left[0,\mathbb{I}\right]$
with $\left[0,\mathbb{I}\right]:=\left\{a\in A|0\leq a\leq\mathbb{I}\right\}$ as candidate for the quantum logic. Note that $0,\mathbb{I}\in L_{A}$ and $p\in L_{A}\Leftrightarrow\mathbb{I}-p\in L_{A}$. Thus we can define the following orthocomplementation on $L_{A}$: $p^{\prime}:=\mathbb{I}-p$. Two elements $p$ and $q$ in $L_{A}$ with $q\leq p^{\prime}$ or equivalently $p+q\leq\mathbb{I}$ are called orthogonal. Note that all non-zero elements in $L_{A}$ carry the norm $1$.
An atom $e$ in this quantum logic is a minimal non-zero element in $L_{A}$; this means: $0\neq q\in L_{A},q\leq e\Rightarrow q=e$. The information capacity $m$ of $A$ or $L_{A}$ is the maximum number of pairwise orthogonal non-zero elements in $L_{A}$. Since such elements are linearly independet, $m$ is finite if the dimension of $A$ is finite, but the dimension of $A$ needs not be finite if $m$ is finite.
An order unit space $A$ with finite information capacity $m<\infty$ is here called spectral, if each element $a\in A$ can be written as $a=\sum s_{k}e_{k}$ with atoms $e_{1},...,e_{n}\in L_{A}$, $s_{1},...,s_{n}\in\mathbb{R}$, $n\leq m$, and $\sum e_{k}=\mathbb{I}$. Spectrality ensures that $L_{A}$ becomes sufficiently large; without it $L_{A}$ might contain only the two elements 00 and $\mathbb{I}$ (although $A\neq\mathbb{R}\mathbb{I}$). In this form, the spectral representation is not unique and therefore not sufficient to define the probability distribution of the observable $a$; other unique forms are available, but require additional assumptions which are not needed and not discussed here.
A state is a non-negative linear functional $\mu$ on $A$ with $\mu(\mathbb{I})=1$. The state space is denoted by $S_{A}$; it becomes a $w^{*}$-compact convex set in the dual of $A$. The state space is called strongly order determining if
$\left\{\mu\in S_{A}|\mu(p)=1\right\}\subseteq\left\{\mu\in S_{A}|\mu(q)=1\right\}\Rightarrow p\leq q$
holds for the $p,q\in L_{A}$.
In any order unit space there is a state $\mu$ with $\mu(a)=\left\|a\right\|$ for each $a\in A$. Particularly for each $e_{j}$ in the above spectral resolution of $a\in A$ there is a state $\mu_{j}$ with $\mu_{j}(e_{j})=1$; then $\mu_{j}(e_{k})=0$ for $k\neq j$ and $\mu_{j}(a)=s_{j}$. Therefore $s_{k}\geq 0$ for each $k$, if $a\geq 0$, and $0\leq s_{k}\leq 1$ for each $k$, if $0\leq a\leq\mathbb{I}$.
3 First findings
We shall now see that $L_{A}=ext\left[0,\mathbb{I}\right]$ becomes a convenient quantum logic, when the order unit space $A$ is spectral.
Lemma 1: If $A$ is a spectral order unit space with finite information capacity $m$ and strongly order determining state space $S_{A}$, then $L_{A}$ is an orthomodular lattice.
Proof. We first show that $p-q$ lies in $L_{A}$, when $q\leq p$ holds with any $p,q\in L_{A}$. We have to verify that $p-q$ is an extreme point of $\left[0,\mathbb{I}\right]$. Thus suppose $p-q=sa+(1-s)b$ with $10<s<$ 1 and $a,b\in\left[0,\mathbb{I}\right]$. For any state $\mu$ with $\mu(q)=1$ we get $\mu(p)=1$ and $0=\mu(p-q)=s\mu(a)+(1-s)\mu(b)$, thus $0=\mu(a)=\mu(b)$ and $1=\mu(\mathbb{I}-a)=\mu(\mathbb{I}-b)$. Because $S_{A}$ is strongly order determining, we have $q\leq\mathbb{I}-a$ and $q\leq\mathbb{I}-b$ or, equivalently, $a+q\leq\mathbb{I}$ and $b+q\leq\mathbb{I}$. Since $p$ is an extreme point, we get from $p=s(a+q)+(1-s)(b+q)$ that $a+q=p=b+q$. Therefore we have $p-q=a=b$ and we have verified that $p-q$ is an extreme point.
Now suppose $p$ and $q$ are orthogonal in $L_{A}$. Then $q\leq\mathbb{I}-p$ and $\mathbb{I}-p-q\in L_{A}$. Hence $p+q=(\mathbb{I}-p-q)^{\prime}\in L_{A}$. Particularly any sum of orthogonal atoms $e_{1}+...+e_{n}$ becomes an element in $L_{A}$.
If $q\in L_{A}$ is not an atom, there is an atom $e_{1}$ with $e_{1}\leq q$. If $q-e_{1}\in L_{A}$ is not an atom, there is a further atom $e_{2}$ with $e_{2}\leq q-e_{1}$. Continuing this process, we finally get $q=e_{1}+...+e_{n}$ with certain atoms $e_{1},...,e_{n}$, since this process stops after at most $m$ steps, where $m$ is the information capacity. Any element in $L_{A}$ thus becomes the sum of atoms.
We now consider two elements $p,q\in L_{A}$ and a spectral representation of their sum $p+q=\sum s_{k}e_{k}$; $e_{1},...,e_{n}$ are atoms with $\sum e_{k}=\mathbb{I}$. Then $0\leq s_{k}\leq 2$ for each $k$. We first assume $s_{k}<2$ for all $k$. If $0\neq y\in L_{A}$ with $y\leq p$ and $y\leq q$, there is $\mu\in S_{A}$ with $\mu(y)=1$. We have $2y\leq p+q$. Thus we get the contradiction $22\leq\mu(p+q)=s_{1}\mu(e_{1})+...+s_{n}\mu(e_{n})\leq max\left\{s_{1},...,s_{n}\right\}<$ 2. Therefore we have $p\wedge q=0$. We now assume $s_{k}=2$ for some $k$ and define $x:=\Sigma_{k:s_{k}=2}\ e_{k}\in L_{A}$. Suppose $\mu(x)=1$ with $\mu\in S_{A}$. The spectral representation then gives $\mu(p+q)=2$. Since $\mu(p)\leq 1$ and $\mu(q)\leq 1$, we get $\mu(p)=1=\mu(q)$. Therefore $x\leq p$ and $x\leq q$ because $S_{A}$ is strongly order determining. Now suppose $y\in L_{A}$ with $y\leq p$ and $y\leq q$. For any $\mu\in S_{A}$ with $\mu(y)=1$ we get $\mu(p)=1=\mu(q)$ and $\mu(p+q)=2$. The spectral representation gives $\mu(x)=1$. Therefore $y\leq x$ and we finally have $x=p\wedge q$. The supremum $p\vee q$ exists in $L_{A}$ since it is identical with $(p^{\prime}\wedge q^{\prime})^{\prime}$ (see [8] for a very similar proof of the lattice properties).
Suppose again that $p$ and $q$ are orthogonal in $L_{A}$. Then $p=e_{1}+...+e_{k}$ and $q=e_{k+1}+...+e_{l}$ with atoms $e_{1},...,e_{l}$. Moreover $\mathbb{I}-e_{1}-...-e_{l}=e_{l+1}+...+e_{n}$ with $n\leq m$. Then $p^{\prime}+q^{\prime}=e_{1}+...+e_{l}+2(e_{l+1}+...+e_{n})$ and $p^{\prime}\wedge q^{\prime}=e_{l+1}+...+e_{n}$ (see above). Thus $p\vee q=(p^{\prime}\wedge q^{\prime})^{\prime}=e_{1}+...+e_{l}=p+q$.
Now we consider $p,q\in L_{A}$ with $q\leq p$ and show $p-q=p\wedge q^{\prime}$. We have $p-q=p\in L_{A}$. Obviously $p-q\leq p$, $p-q\leq\mathbb{I}-q$ and therefore $p\wedge q^{\prime}\leq p-q$. If $x\in L_{A}$ with $p\leq x$ and $q^{\prime}\leq x$, then $p-q\leq p\leq x$ and $p-q\leq\mathbb{I}-q=q^{\prime}\leq x$ and therefore $p-q=p\wedge q^{\prime}$.
Finally $p=q+(p-q)=q+p\wedge q^{\prime}=q\vee(p\wedge q^{\prime})$ and we have verified the orthomodular law. $\square$
Note that, with the finite information capacity, the orthomodular lattice $L_{A}$ becomes atomic and complete.
With an order unit space $A$ and any $p\in L_{A}$, let $A_{p}$ denote the set of those elements $a\in A$ for which there is a non-negative real number $s$ with $-sp\leq a\leq sp$; $A_{p}$ is an order unit space with the order unit $p$.
Lemma 2: Let $A$ be a spectral order unit space with finite information capacity $m$ and a strongly order determining state space $S_{A}$.
Each $a\in A_{p}$ with $p\in L_{A}$ has a spectral representation $a=\sum s_{k}e_{k}$ with $e_{k}\leq p$ for each $k$ with $s_{k}\neq 0$.
Moreover, $L_{A_{p}}=ext(\left[0,p\right])$ coincides with $\left\{q\in L_{A}|q\leq p\right\}$ and $A_{p}$ is spectral.
Proof. Assume $0\leq a\leq rp$ with some $p\in L_{A}$ and $0<r\in\mathbb{R}$ and consider the spectral representation $a=\sum s_{k}e_{k}$, $0\leq s_{k}$ for each $k$ and the $e_{k}$ are atoms with $\sum e_{k}=\mathbb{I}$. Let $\mu\in S_{A}$ be any state with $\mu(p^{\prime})=1$. Then $\mu(p)=0$ and $0=\mu(a)=\sum s_{k}\mu(e_{k})$. Therefore $\mu(e_{k})=0$ and $\mu(e_{k}^{\prime})=1$ for each $k$ with $s_{k}\neq 0$. Since $S_{A}$ is strongly order determining, we thus have $p^{\prime}\leq e_{k}^{\prime}$ and $e_{k}\leq p$ for each $k$ with $s_{k}\neq 0$.
If $-rp\leq a\leq rp$, then $a+rp=\sum s_{k}e_{k}$ with $e_{k}\leq p$ for $s_{k}\neq 0$. With $p_{o}:=\sum_{s_{k}\neq 0}e_{k}$ we have $p_{o}\leq p$ and $a=\sum s_{k}e_{k}-rp=\sum s_{k}e_{k}-rp_{o}-r(p-p_{o})=\sum_{s_{k}\neq 0}(s_{k}-r)e_{k}-r(p-p_{o})$. Represent $p-p_{o}$ as a sum of atoms to complete the proof of the first part of the lemma.
Now suppose $x\in ext(\left[0,p\right])$. Then $0\leq x\leq p$ and $x=\sum s_{j}e_{j}$ with $0<s_{j}$ and $e_{j}\leq p$. If $s_{j_{o}}\neq 1$ for some $j_{o}$, we have $x=s_{j_{o}}(e_{j_{o}}+\sum_{j\neq j_{o}}s_{j}e_{j})+$ $(1-s_{j_{o}})(\sum_{j\neq j_{o}}s_{j}e_{j})$ and $x$ is not extremal. Thus we have $x=\sum e_{j}\in L_{A}$ Vive versa, if $x\in L_{A}$ with $x\leq p$, $x$ becomes extremal in $\left[0,p\right]$. $\square$
The state space $S_{A_{p}}$ consists of the restrictions to $A_{p}$ of the states $\mu\in S_{A}$ with $\mu(p)=1$ and is strongly order determining.
Two elements $p_{1}$ and $p_{2}$ in any orthomodular lattice $L$ are called compatible, if $L$ contains three orthogonal elements $q_{1},q_{2},q_{3}$ with $p_{1}=q_{1}+q_{2}$ and $p_{2}=q_{2}+q_{3}$; then $q_{2}=p_{1}\wedge p_{2}$. The center is the set of all those elements that are compatible with each element in $L$, and $L$ is called irreducible if the center is trivial, which means that the center is $\left\{0,\mathbb{I}\right\}$. Generally, the center is a Boolean sublattice and its minimal non-zero elements are pairwise orthogonal. Note that these elements are minimal in the center, but not in $L$. Each atom in $L$ lies under one of the central minimal non-zero elements.
Lemma 3: Suppose $A$ is a spectral order unit space with finite information capacity $m$ and a strongly order determining state space $S_{A}$. Let $c_{1},...c_{n}$ be the minimal non-zero elements in the center of $L_{A}$. Then $n\leq m$ and $A=A_{c_{1}}\oplus A_{c_{2}}\oplus...\oplus A_{c_{n}}$.
Proof. Each $a\in A$ has a spectral representation $a=\sum s_{k}e_{k}$ with atoms $e_{1},...,e_{n}\in L_{A}$, $s_{1},...,s_{n}\in\mathbb{R}$ and $\sum e_{k}=\mathbb{I}$. For each $e_{k}$ there is a $c_{j}$ with $e_{k}\leq c_{j}$. Therefore we have $a\in A_{c_{1}}\oplus A_{c_{2}}\oplus...\oplus A_{c_{n}}$. $\square$
4 The gbit property
Spaces with the information capacity $m=2$ can be considered generalized models of a bit. We would expect that any pair of atoms generates such a model. This motivates the following postulate for a spectral order unit space $A$ with finite information capacity and strongly order determiningg state space:
For any two atoms $e_{1}\neq e_{2}$, $A_{e_{1}\vee e_{2}}$ has the information capacity m = 2.
This condition shall here be called gbit-property. It can be stated in the following equivalent way:
If $e_{1}$ and $e_{2}$ are atoms and $0\neq q\leq e_{1}\vee e_{2}$ holds with $q\in L_{A}$, then either $q=e_{1}\vee e_{2}$ or $q$ is an atom.
We shall now see that the gbit property is nothing else as the covering property in disguise. The covering property was studied a lot in lattice theory [1, 2, 22] and looks as follows:
If $e$ is an atom and $p\leq q\leq p\vee e$ for any $p,q$ in the lattice, then $q=p$ or $q=p\vee e$.
If the covering property holds in an orthomodular lattice, there is a dimension function $dim$ such that each element $p$ is the supremum of $dim(p)$ orthogonal elements in $L$ [1].
Lemma 4: The following two conditions are equivalent for any orthomodular lattice $L$:
  • (i)
    If $0\neq q\leq e_{1}\vee e_{2}$ with $q,e_{1},e_{2}\in L$, where $e_{1}$ and $e_{2}$ are atoms, then either $q=e_{1}\vee e_{2}$ or $q$ is an atom (gbit-property).
  • (ii)
    $L$ has the covering property.
Proof. (i) $\Rightarrow$ (ii): The proof becomes identical with the second part of the proof of [2] Proposition 9.11. The lattice of projective units is considered there, but the line of reasoning works for any orthomodular lattice.
(ii) $\Rightarrow$ (i): Suppose $L$ has the covering property and assume $0\neq q\leq e_{1}\vee e_{2}$ with $q,e_{1},e_{2}\in L$, where $e_{1}$ and $e_{2}$ are atoms. Then $1=dim(q)$ and $q$ is an atom, or $22\leq dim(q)\leq dim(e_{1}\vee e_{2})\leq dim(e_{1})+dim(e_{2})=$ 2; in the second case we have $dim(q)=dim(e_{1}\vee e_{2})$ and then $q=e_{1}\vee e_{2}$. $\square$
5 Continuous symmetry
An automorphism of an order unit space $A$ is a an invertible positive transformation $T:A\rightarrow A$ with a positive inverse and $T(\mathbb{I})=\mathbb{I}$. The automorphisms form a group denoted by $Aut(A)$. This becomes a compact group, if the dimension of $A$ is finite. A spectral order unit space $A$ shall here be called continuously symmetric, if the following condition holds:
For any pair of atoms $e_{1}$ and $e_{2}$ there is a continuous map $t\rightarrow T_{t}$ from the real unit interval $\left[0,1\right]$ to $Aut(A)$ such that $T_{0}$ is the identity in $Aut(A)$ and $T_{1}(e_{1})=e_{2}$.
This means that any atom $e_{1}$ can be continuously transformed to any other atom $e_{2}$ in a reversible way and is one of Hardy’s five axioms [11], although he assumes this for the pure states instead of atoms. In some cases the following weaker non-continuous form of symmetry will be sufficient; $A$ shall be called weakly symmetric, if this holds:
For any pair of atoms $e_{1}$ and $e_{2}$ there is $T\in Aut(A)$ with $T(e_{1})=e_{2}$.
Weak symmetry means that the automorphism group $Aut(A)$ acts transitively on the atoms. Before we come to our main result, we prove two auxiliary lemmas.
Lemma 5: Suppose that $A$ is a spectral order unit space with the finite information capacity $m$ and a strongly order determining state space. If $A$ is continuously symmetric, then $A$ is irreducible.
Proof. Suppose that $L_{A}$ is not irreducible. Then there are at least two minimal non-zero elements $c_{1}$ and $c_{2}$ in the center and we select atoms $e_{1}$ and $e_{2}$ in $L_{A}$ such that $e_{1}\leq c_{1}$ and $e_{2}\leq c_{2}$ Let $T_{t}\in Aut(A)$, $t\in\left[0,1\right]$, be a continuous transformation with $T_{0}(e_{1})=e_{1}$ and $T_{1}(e_{1})=e_{2}$. For each $t$, $T_{t}(e_{1})$ is an atom. Define $s:=inf\left\{t:T_{t}(e_{1})\leq c_{2}\right\}$. The continuity then implies $T_{s}(e_{1})\leq c_{2}$ and thus $0<s$. Note that the set $\left\{a\in A|0\leq a\leq p\right\}$ is closed for any $p\in L_{A}$. For $t<s$ we have $T_{t}(e_{1})\leq\mathbb{I}-c_{2}$ and the continuity implies $T_{s}(e_{1})\leq\mathbb{I}-c_{2}$. We then have the contradiction $00\neq T_{s}(e_{1})\leq c_{2}\wedge c^{{}^{\prime}}_{2}=$ 0. $\square$
Lemma 6: Let $A$ be a finite-dimensional spectral order unit space with a strongly order determining state space. If $A$ is weakly symmetric, then $A$ possesses an $Aut(A)$-invariant inner product $\left\langle\ |\ \right\rangle_{o}$ with $\left\langle e|e\right\rangle_{o}=1$ for each atom $e$.
[…]
for $a,b\in A$. Since $Aut(A)$ acts transitively on the atoms, we have $0\neq\left\langle p|p\right\rangle_{o}=\left\langle q|q\right\rangle_{o}$ for any two atoms $p$ and $q$ and we can normalize $\left\langle\ |\ \right\rangle_{o}$ in such a way that $\left\langle e|e\right\rangle_{o}=1$ for every atom $e$. $\square$
Theorem 1: Suppose that $A$ is a spectral order unit space with finite dimension, information capacity $m$ and a strongly order determining state space.
If $m=2$ and $A$ is weakly symmetric, then $A$ is a spin factor (the spin factors are the irreducible Euclidean Jordan algebras with $m=2$).
If $m\geq 4$, continuous symmetry and the gbit-property are given, then $A$ is a simple (irreducible) Euclidean Jordan algebra and can be represented as the Jordan matrix algebra of Hermitian $m\times m$-matrices over the real numbers, complex numbers or quaternions.
Proof. (i) Suppose $m=2$ and $A$ is weakly symmetric. If $e_{1}$ and $e_{2}$ are atoms, there is $T\in Aut(A)$ with $T(e_{1})=e_{2}$. Then $T(e_{1}^{\prime})=e_{2}^{\prime}$ and $\left\langle e_{1}|e_{1}^{\prime}\right\rangle_{o}=\left\langle T(e_{1})|T(e_{1}^{\prime})\right\rangle_{o}=\left\langle e_{2}|e_{2}^{\prime}\right\rangle_{o}$ with the $Aut(A)$-invariant inner product from Lemma 6. So we get a real number $s_{o}$ with $\left\langle e,e^{\prime}\right\rangle_{o}=s_{o}$ for all atoms $e$.
With the Cauchy-Schwarz inequality we get $\left|\left\langle e,e^{\prime}\right\rangle_{o}\right|\leq\left\langle e,e\right\rangle_{o}\left\langle e^{\prime},e^{\prime}\right\rangle_{o}=1$ and $\left|\left\langle e,e^{\prime}\right\rangle_{o}\right|=1$ is impossible, since $e$ and $e^{\prime}$ are linearly independent. Thus $\left|s_{o}\right|<1$, what we now use to define a further inner product on $A$ (see [18] for a similar construction):
$\left\langle a|b\right\rangle_{1}:=\frac{1}{1-s_{o}}\left[\left\langle a|b\right\rangle_{o}-\frac{s_{o}}{(1+s_{o})^{2}}\left\langle\mathbb{I}|a\right\rangle_{o}\left\langle\mathbb{I}|b\right\rangle_{o}\right]$
for $a,b\in A$. For the atoms $e$ we have $\left\langle\mathbb{I}|e\right\rangle_{o}=\left\langle e+e^{\prime}|e\right\rangle_{o}=1+s_{o}$ and therefore $\left\langle e|e\right\rangle_{1}=1$ and $\left\langle e|e^{\prime}\right\rangle_{1}=0$. Furthermore, any $a\in A$ has the form $a=se+te^{\prime}$ with an atom $e$ and $s,t\in\mathbb{R}$ and $\left\langle a|a\right\rangle_{1}=\left\langle se+te^{\prime}|se+te^{\prime}\right\rangle_{1}=s^{2}+t^{2}$. Thus $\left\langle a|a\right\rangle_{1}\geq 0$ and $\left\langle a|a\right\rangle_{1}=0$ iff $a=0$.
With $V:=\left\{v\in A|\left\langle\mathbb{I}|v\right\rangle_{1}=0\right\}$ we have $A=\mathbb{R}\mathbb{I}\oplus V$. For $v,w\in V$ and $s,t\in\mathbb{R}$ we define the following bilinear commutative product on $A$:
$(v+s\mathbb{I})\circ(w+t\mathbb{I})=tv+sw+(\frac{1}{4}\left\langle v|w\right\rangle_{1}+st)\mathbb{I}.$
With this product $A$ becomes a spin factor ([2] 3.33) and spin factors are Jordan algebras ([2] 3.37). Note that any positive multiple of an inner product provides another one and the parameter $1/4$ is conveniently chosen here to make sure that the atoms become idempotent elements, what we will show now.
Obviously $\mathbb{I}$ is the identity for this product. If $e$ is an atom, then $e^{\prime}$ is an atom because of $m=2$ and $\left\langle\mathbb{I}|e-e^{\prime}\right\rangle_{1}=\left\langle e+e^{\prime}|e-e^{\prime}\right\rangle_{1}=\left\langle e|e\right\rangle_{1}-\left\langle e^{\prime}|e^{\prime}\right\rangle_{1}=0$. This means $e-e^{\prime}\in V$. Furthermore $\left\langle e-e^{\prime}|e-e^{\prime}\right\rangle_{1}=2$, $e=(e-e^{\prime})/2+\mathbb{I}/2$ and $e^{2}=e\circ e=(e-e^{\prime})/2+(\left\langle e-e^{\prime}|e-e^{\prime}\right\rangle_{1}/8+1/4)\ \mathbb{I}=(e-e^{\prime})/2+1/4\mathbb{I}=e$.
Moreover $e\circ e^{\prime}=e\circ(\mathbb{I}-e)=e-e^{2}=0$ and $(se+te^{\prime})^{n}=s^{n}e+t^{n}e^{\prime}$ for any $s,t\in\mathbb{R}$, $n\in\mathbb{N}$. Thus we have $a^{2}\geq 0$ for any $a\in A$, $a^{2}=0$ iff $a=0$ and each positive element in $A$ is the square of some other element.
(ii) Suppose $m\geq 4$ and that the continuous symmetry and the gbit-property are given. We know from Lemma 1 that $L_{A}=ext\left[0,\mathbb{I}\right]$ is an orthomodular lattice, which becomes atomic and complete since the information capacity is finite. By Lemma 4, $L_{A}$ possesses the covering property and by Lemma 5, $L_{A}$ is irreducible.
Every irreducible complete orthomodular lattice with the covering property and $m\geq 4$ is isomorphic to the lattice of orthoclosed subspaces of a m-dimensional inner product space over some a division ring [22]. This division ring can be rather exotic [12]. MacLaren [14] identified the following two conditions that make the division ring become the real numbers, complex numbers or quaternions:
  • (1)
    The set of all atoms $e$ with $e\leq p$ is compact for any $0\neq p\in L_{A}$.
  • (2)
    For some $p\in L_{A}$ and real interval $I$, there exists a continuous nonconstant function $t\rightarrow q_{t}$ from $I$ to $\left\{q\in L_{A}|q\leq p\right\}$.
We first show (1). Select any atom $e_{o}$. The set of all atoms is identical with $\left\{T(e_{o})|T\in Aut(A)\right\}$ and becomes compact therefore. For any $0\neq p\in L_{A}$, the intersection of this set with the closed set $\left\{a\in A|0\leq a\leq p\right\}$ is then compact as well and we have (1).
Condition (2) is fulfilled with $p:=\mathbb{I}$, $I:=\left[0,1\right]$, any atoms $e_{1}\neq e_{2}$ and $q_{t}:=T_{t}(e_{1})$, where the $T_{t}$ are the automorphisms that transform $e_{1}$ continuously to $e_{2}$.
Hence $L_{A}$ is isomorphic to the lattice of closed subspaces of the m-dimensional Hilbert space over the real numbers, complex numbers or quaternions. This lattice is isomorphic to the projection lattice in the Jordan matrix algebra over the division ring and $A$ becomes isomorphic to this algebra. $\square$
MacLaren’s result [14] bases upon earlier work [13, 24, 25, 26] and ultimately on Pontrjagin’s theorem which states that there are only three (topologically) connected division rings, namely the real numbers, complex numbers and the quaternions [23].
The proof of the above theorem is extremely different in the two cases with information capacity $m=2$ and $m\geq 4$. The weak symmetry is sufficient in the first case and the gbit-property follows directly from $m=2$. The case $m=3$ remains a little enigmatic. Piron’s representation theorem for the orthomodular lattices [22] is not applicable then. The Hermitian $3\times 3$-matrices over the real numbers, complex numbers, quaternions and in this case also over the octonions form Euclidian Jordan algebras with $m=3$ and satisfy all our postulates, but the existence of further irreducible examples with $m=3$ remains unclear here.
6 Conclusions
Any physical theory needs mathematical models for the dynamical processes. When we assume that these processes are reversible and continuous in time and that the atoms can continuously evolve from one to each other, we come to the symmetry postulate that we have studied here. It rules out the classical and the reducible finite-dimensional cases and brings us to the simple (irreducible) Euclidean Jordan algebras (spin factors and Jordan matrix algebras over the real numbers, complex numbers and quaternions). It thus becomes one of the basic principles from which quantum theory could be reconstructed.
The question why nature in quantum theory prefers the complex numbers over the real numbers and the quaternions is not considered here, but has already been addressed in different ways. One way is to consider locally tomographic models for composite systems [4, 5, 6, 7, 15, 20]. Other ways are to postulate somehow that an observable plays the role of the Hamiltonian (dynamcial correspondence [2, 19], energy observability [3]).
The case with information capacity $m=3$ is somewhat exceptional and remains a little enigmatic. Furthermore, many examples are known with $m=2$ and infinite dimension (for instance the spin factors with infinite dimension), but it is not evident whether the case $3\leq m<\infty$ includes infinite-dimensional order unit spaces. We had to presuppose the finite dimension to get a compact automorphism group.
We have needed the gbit-property, which turned out to be the covering property in disguise. The physical or information theoretical necessity of these properties might be disputable and it is not clear what mathematical structure would arise when we do without them. We have not needed any one of the following postulates that are part of various reconstructions of quantum theory: symmetry of the transition probability [21], bit symmetry [18], strong symmetry [3], absence of third-order interference [3, 21] and the existence of compressions or filters [2, 9, 10, 16]. These postulates are satisfied by the Jordan matrix algebras and thus become deducible features in our approach.
The compressions are an abstract mathematical notion with an important physical meaning; they represent an axiomization of the quantum measurement process [2]. In the approach presented here their existence is a consequence of the continuous symmetry.